From a box containing 5 dimes and 3 nickels, 4 coins are selected at random without replacement.

Find the probability distribution for the total T of the 4 coins. Express the probability distribution graphically as a probability histogram. Find the Expected Value and Standard Deviation of the number of coins.


1
Expert's answer
2021-08-08T17:00:49-0400
"P(4D \& 0N)=\dfrac{\dbinom{5}{4}\dbinom{3}{0}}{\dbinom{5+3}{4}}=\dfrac{1}{14}"

"P(3D \& 1N)=\dfrac{\dbinom{5}{3}\dbinom{3}{1}}{\dbinom{5+3}{4}}=\dfrac{3}{7}"

"P(2D \& 2N)=\dfrac{\dbinom{5}{2}\dbinom{3}{2}}{\dbinom{5+3}{4}}=\dfrac{3}{7}"

"P(1D \& 3N)=\dfrac{\dbinom{5}{1}\dbinom{3}{3}}{\dbinom{5+3}{4}}=\dfrac{1}{14}"


"\begin{matrix}
 T & 0.40 & 0.35 & 0.30 & 0.25\\
\\
 p(T) & \dfrac{1}{14} & \dfrac{3}{7} & \dfrac{3}{7}& \dfrac{1}{14}
\end{matrix}"

"E(T)=\dfrac{1}{14}(0.40)+\dfrac{3}{7}(0.35)+\dfrac{3}{7}(0.30)+\dfrac{1}{14}(0.25)"

"=0.325"

"E(T^2)=\dfrac{1}{14}(0.40)^2+\dfrac{3}{7}(0.35)^2+\dfrac{3}{7}(0.30)^2"

"+\dfrac{1}{14}(0.25)^2=\dfrac{1.4975}{14}"

"Var(T)=\sigma^2=E(T^2)-(E(T))^2"

"=\dfrac{1.4975}{14}-(0.325)^2\approx0.001339"

"\sigma=\sqrt{\sigma^2}\approx\sqrt{0.001339}\approx0.0366"